OpenStudy (anonymous):

write equation in slope-intercept form 6x-4y=8

OpenStudy (kira_yamato):

4y = 6x - 8 \[y = \frac{3}{2}x-2\]

OpenStudy (anonymous):

Slope intercept is when the equation is modeled after the general form y=mx+b Where m is the slope of the line and b is the y-intercept Doing this is fairly simple, you need only to isolate the y variable on the left side of the equation. -4y=8-6x (subtract 6x) y=(3x)/2 - 2 (divide by -4) And there you have your slope-intercept equation.

OpenStudy (anonymous):

so is Kira wrong

OpenStudy (kira_yamato):

It's the same, because he just brought the x into the numerator, while I left the x out of the numerator

OpenStudy (anonymous):

No, we have the same equation, he just knows how to use the equation maker.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

how do i write this in standard form, 3y=3/4x+5 the "a" must be positive and non fraction

OpenStudy (kira_yamato):

Do you mean \[3y = \frac{3}{4}x + 5\]?